Q#2 A machine acquired at the beginning of the fiscal year at a cost of Rs. 45,200 has an estimated residual value of Rs. 2,000 and an estimated useful life of 8 years. Determine the depreciation for each of the first five years 

(a) by the straight line method, 

(b) by the declining-balance method, using twice the straight line rate (25%) and 

(c) by the sum of the year digit method.


1
Expert's answer
2021-04-23T12:07:40-0400

a)


"\frac{45200-2000}{8}=5400"

b)


"0.25\times (45200-2000)=10800"


"0.25\times (45200-2000-10800)=8100"

"0.25\times(45200-2000-10800-8100)=6075"

"0.25\times(45200-2000-10800-8100-6075)=4556.25"

"0.25\times 13668.75=3417"

c)


"\frac {45200\times8}{36}=10044"

"\frac {45200\times7}{36}=8788"

"\frac {45200\times6}{36}=7533"

"\frac {45200\times5}{36}=6277"


"\frac {45200\times4}{36}=5022"